OpenStudy (anonymous):

the width of a rectangle is 5 inches less than its length. its area is 50square inches. find the deminsions of the rectangle

OpenStudy (anonymous):

area of a rectangle A = W * h in this case: W = h-5 A = (h-5) * h = 50 -> h^2 - 5h = 50 h^2 - 5h - 50 = 0 h= 10 W = 10-5 = 5 height: 10 inches width: 5 inches

OpenStudy (anonymous):

thank you